# Neural Machine Translation of Clinical Text: An Empirical Investigation into Multilingual Pre-Trained Language Models and Transfer-Learning

### 12 Dec 2023 v1

@@ Abstract
Clinical text and documents contain very rich information and knowledge in healthcare, and their processing using state-of-the-art language technology becomes very important for building intelligent systems for supporting healthcare and social good. This processing includes creating language understanding models and translating resources into other natural languages to share domain-specific cross-lingual knowledge. In this work, we conduct investigations on clinical text machine translation by examining multilingual neural network models using deep learning such as Transformer based structures. Furthermore, to address the language resource imbalance issue, we also carry out experiments using a transfer learning methodology based on massive multilingual pre-trained language models (MMPLMs). The experimental results on three sub-tasks including 1) clinical case (CC), 2) clinical terminology (CT), and 3) ontological concept (OC) show that our models achieved top-level performances in the ClinSpEn-2022 shared task on English-Spanish clinical domain data. Furthermore, our expert-based human evaluations demonstrate that the small-sized pre-trained language model (PLM) won over the other two extra-large language models by a large margin, in the clinical domain fine-tuning, which finding was never reported in the field. Finally, the transfer learning method works well in our experimental setting using the WMT21fb model to accommodate a new language space Spanish that was not seen at the pre-training stage within WMT21fb itself, which deserves more exploitation for clinical knowledge transformation, e.g. to investigate into more languages. These research findings can shed some light on domain-specific machine translation development, especially in clinical and healthcare fields. Further research projects can be carried out based on our work to improve healthcare text analytics and knowledge transformation.
@@ 1 Introduction
Healthcare Text Analytics (HECTA) have gained more attention nowadays from researchers across different disciplines, due to their impact on clinical treatment, decision-making, hospital operation, and their recently empowered capabilities. These developments have much to do with the latest development of powerful language models (LMs), advanced machine-learning (ML) technologies, and increasingly available digital healthcare data from social media [Griciūtė et al., 2023, Oyebode et al., 2021, Luo et al., 2022] and discharged outpatient letters from hospital settings [Henry et al., 2020, Spasic et al., 2020, Percha, 2021].
@@
Intelligent healthcare systems have been deployed in some hospitals to support clinicians’ diagnoses and decision-making regarding patients and problems [Noor et al., 2022, Qian et al., 2021]. Such usages include key information extraction (IE) from electronic health records (EHRs), normalisation to medical terminologies, knowledge graph (KG) construction, and relation extraction (RE) between symptoms (problems), diagnoses, treatments, and adverse drug events [Wu et al., 2022, Nguyen et al., 2023]. Some of these digital healthcare systems can also help patients self-diagnose in situations where no General Practitioners (GPs) and professional doctors are available [Wroge et al., 2018, Zhu et al., 2021].
@@
However, due to the language barriers and inequality of digital resources across languages, there is an urgent need for knowledge transformation, such as from one human language to another [Costa-jussà et al., 2022, Khoong and Rodriguez, 2022]. Thus, to help address digital health inequalities, machine translation (MT) technologies can be of good use in this case.
@@
MT is one of the earliest artificial intelligence (AI) branches dating back to the 1950s, and it has gained a boom with other natural language processing (NLP) tasks in recent years due to the newly designed powerful learning model Transformers [Weaver, 1955, Vaswani et al., 2017, Devlin et al., 2018, Han et al., 2021]. Several attention mechanisms designed in Transformer deep neural models are proven to be capable of better learning from a large amount of available digital data compared to traditional statistical and neural network-based models [Han, 2022, Kuang et al., 2018, Han and Kuang, 2018].
@@
In this work, we investigate the state-of-the-art Transformer based Neural MT (NMT) models regarding clinical domain text translation, to facilitate digital healthcare and knowledge transformation with the workflow drawn in Figure 1. Being aware of some current development in the competition of language model sizes in NLP field, we set up the following base models for comparison study: 1) a small-sized multilingual pre-trained language model (s-MPLM) Marian, which was developed by researchers at the Adam Mickiewicz University in Poznan and at the NLP group in University of Edinburgh [Junczys-Dowmunt et al., 2018a, Junczys-Dowmunt et al., 2018b]; and 2) a massive-sized multilingual pre-trained LM (MMPLM/xL-MPLM) NLLB, developed by Meta-AI covering more than 200 languages [Costa-jussà et al., 2022]. In addition to this, we set up a third model to investigate the possibility of transfer learning in the clinical domain MT: 3) the WMT21fb model which is another MMPLM from Meta-AI but with a limited amount of pre-trained language pairs including from English to Czech, German, Hausa, Icelandic, Japanese, Russian, and Chinese, and the opposite [Tran et al., 2021].
@@
The testing language pairs of these translation models in our work are English $\leftrightarrow$ Spanish. There aren’t other language pairs of openly available resources in the clinical domain MT as far as we know. We use the international shared task challenge data from ClinSpEn2022 “clinical domain Spanish-English MT 2022” for this purpose  [^1] https://codalab.lisn.upsaclay.fr/competitions/6696. ClinSpEn2022 was a sub-task of the BioMedical MT track at WMT2022 [Neves et al., 2022]. There are three translation tasks inside ClinSpEn2022 including i) clinical cases report; ii) clinical terms, and 3) ontological concepts from the biomedical domain.
@@
Regarding the evaluation of these LMs, we used the evaluation platform offered by ClinSpEn2022 shared task including several automatic metrics such as BLEU, METEOR, ROUGE, COMET. However, the automatic evaluation results did not give apparent differentiation between models on some tasks. Furthermore, there are issues like in-consistency regarding model ranking across automatic metrics. To address these issues and give a high-quality evaluation, we carried out an expert-based human evaluation step on three models using outputs of Task one “clinical case report”.
@@
Our experimental investigation shows that 1) the extra-large MMPLM does not necessarily win the small-sized MPLM on clinical domain MT via fine-tuning; 2) our transfer-learning model works successfully for clinical domain MT task on language pairs that were not pre-trained upon but using fine-tuning. The first finding can shed some light on the research field that in clinical domain-specific MT, it is worthy to carry out more work on data cleaning and fine-tuning rather than building extra large LMs. Our second finding tells us the capability of MMPLMs in generating a new language pair knowledge space for translating clinical domain text even though this language pair was unseen in the pre-training stage with our experimental settings. This can be useful to low-resource NLP, such as the work by [Almansor and Al-Ani, 2018, Islam et al., 2021].  [^2] This paper reports systematic investigation findings based upon the preliminary work from [Han et al., 2022a, Han et al., 2022b].
@@
The rest of this article is organised as below: Section 2 surveys the related work to ours including clinical domain MT and NLP, large LMs, and transfer learning. Section 3 details the three LMs we deployed for comparison study. Section 4 introduces the experimental work we carried out and automatic evaluation outcomes. Section 5 follows up with expert-based human evaluation and the results. Finally, Section 6 concludes our work with discussion.
@@ 2 Related Work
Applying NLP models to clinical healthcare has attracted much attention of many researchers, such as the work on disease status prediction using discharged summaries by ?), temporal expressions and events extraction from clinical narratives using combined methods of rules and machine learning by ?), using knowledge-based and data-driven methods for de-identification task in clinical narratives by ?), systematic reviews on clinical text mining and healthcare by ?) and ?), etc.
@@
However, using MT to help translating clinical text for knowledge transformation and help clinical decision-making is still a rising topic [Khoong and Rodriguez, 2022], even though it has been proven to be useful in the history for assisting health communication especially with post-editing strategies [Dew et al., 2018]. This is partial because of the sensitive domain and high risk in clinical settings [Randhawa et al., 2013]. Some of the recent progress on using MT for clinical text includes the work by ?) which leverages SNOMED-CT terms [Donnelly and others, 2006] and relations for MT between Basque and Spanish languages, ?) which applies NMT model to identify semantic concepts in “abundant interchangeable words” in clinical domain and their experimental result shows NMT model can greatly improve the efficiency on extracting UMLS [Bodenreider, 2004] concepts from a single document by using 30 milliseconds in comparison to traditional regular expression based methods which takes 3 seconds, and ?) which uses NMT to simplify the typical multi-stage workflow on clinical report dictation and even correct the errors from speech recognition.
@@
With the prevalence of multilingual PLMs (MPLMs) developed from NLP fields, it becomes a current need to test their performances in the clinical domain of NMT. MPLMs have been adopted by many NLP tasks since the first emergence of the Transformer based learning structure [Vaswani et al., 2017]. Among these, Marian is a small-sized MPLM led by Microsoft Translator based upon Nenatus NMT [Sennrich et al., 2017] with around 7.6 million parameters [Junczys-Dowmunt et al., 2018a]. Then, different research and development teams have been competing on the size of their LMs in recent years, e.g. the massive MPLMs (MMPLMs) WMT21fb and NLLB by Meta-AI which have the number of parameters set of 4.7 billion and 54 billion respectively [Tran et al., 2021, Costa-jussà et al., 2022]. To investigate the performances of these different models with varied model sizes towards clinical domain NMT with fine-tuning, we set up all these three models as our base models. To the best of our knowledge, our work is the first one regarding the comparison between small-size and extra-large MPLMs in the clinical domain of NMT.
@@
Very close to the clinical domain, there has been a biomedical domain MT challenge series together held with the Annual Conference of MT (WMT), since 2016 [Bojar et al., 2016, Yeganova et al., 2021]. The historical biomedical MT task covered corpus of biomedical terminologies, scientific abstracts from Medline, summaries of proposals for animal experiments, etc. In 2022, it was the first time that this Biomedical-MT shared task introduced clinical domain data for Spanish-English language pairs [Neves et al., 2022].
@@
As the WMT21fb model does not include Spanish in its pre-training, we also examine the transfer learning technology into the clinical domain NMT towards Spanish-English using the WMT21fb model. Transfer-learning [Alyafeai et al., 2020] has proved useful for text classification and relation extraction [Pomares-Quimbaya et al., 2021, Peng et al., 2019], and low-resource MT [Jiang et al., 2022] fields. However, to the best of our knowledge, we are the first to test clinical domain NMT via transfer learning using MMPLMs.
@@ 3 Experimental Designs
In this section, we introduce more details about the three MPLMs that we investigate in this work, i.e., Marian [Junczys-Dowmunt et al., 2018a], WMT21fb [Tran et al., 2021], and NLLB [Costa-jussà et al., 2022].
@@
Figure 3: Several Attention based Transformer NMT structure [Vaswani et al., 2017].
@@ 3.1 Multilingual Marian NMT
Firstly, we draw a training diagram of the original Marian model on its pre-training steps in Figure 2 according to [Junczys-Dowmunt et al., 2018a]. The pre-processing step includes tokenisation, true-casing, and Byte-Pair Encoding (BPE) for sub-words. The shallow training is to learn a mid-phase translation model to produce temporary target outputs for back-translation. Then, the back-translation step produces the same amount of input source sentences to enlarge the corpus. The deep-training step first uses four left-to-right models which can be RNN [Sennrich et al., 2017] or Transformer [Vaswani et al., 2017] structures, which is followed by four right-to-left models in the opposite direction. The ensemble-decoding step will generate the n-best hypothesis translations for each source input segment, which will be re-ranked using a re-scoring mechanism. Finally, in Marian NMT, there is an automatic post-editing step integrated before producing the output. This step is also based on an end-to-end neural structure by modelling the set(MT-output, source sentence)$\rightarrow$“post-edited output” as introduced by ?).
@@ 3.2 Extra-Large Multilingual WMT21fb and NLLB
Instead of the optional RNN structure used in the Marian model, both massive-sized multilingual PLMs (MMPLMs) WMT21fb and NLLB adopted Transformer as the main methodology. As shown in Figure 3, Transformer’s main components for encoder include position encoding, Multi-Head Attention, and Feed-Forward Network with layer normalisation at both two steps. The decoder uses Masked Multi-Head Attention to constrain the generation model only taking the already generated text into account.
To increase the model capacity without making the extra-large model too slow for training, inspired by the work from ?), the WMT21fb model included “Sparsely Gated Mixture-of- Expert (MoE)” models into the FFN layer of Transformer, as shown in Figure 4. The MoE model will only pass a sub-set of model parameters into the next level, thus decreasing the computational cost. However, this dropout is done is a random manner.
@@
To further improve on this, the NLLB model designed a Conditional MoE Routing layer inspired by ?) to ask the MoE model to decide which tokens to dropout according to their capacity demanding or routing efficiency. This is achieved by a binary gate, which assigns weights to dense FNN $FFN_{shared}$ or MoE Gating, as in Figure 5. The Conditional MoE also removes language-specific parameters for learning.
In summary, the WMT21fb and NLLB models share very similar learning structures, but most differently WMT21fb used language-specific constrained learning. The WMT21fb model we applied is ‘wmt21-dense-24-wide.En-X’ (and X-En direction) which has 4.7 billion parameters  [^4] https://github.com/facebookresearch/fairseq/tree/main/examples/wmt21 and contains the language pairs English $\leftrightarrow$ Chinese, Czech, German, Hausa, Icelandic, Japanese, and Russian. The full NLLB model includes 200+ languages and has 54.5 billion parameters. Due to the computational restriction, we applied the distilled model of NLLB, i.e. NLLB-distilled, which has 1.3 billion parameters.
The WMT21fb model does not have Spanish in the trained language pairs, while NLLB includes Spanish as a high-resource language. This is a perfect setting for us to examine the transfer-learning technology on the clinical domain NMT by fine-tuning a translation model for the Spanish language on the WMT21fb model and comparing the output with the NLLB model (Spanish version).

### 21 Feb 2024 v2

@@ Abstract
Clinical texts and documents contain a wealth of information and knowledge in the field of healthcare, and their processing, using state-of-the-art language technology, has become very important for building intelligent systems capable of supporting healthcare and providing greater social good. This processing includes creating language understanding models and translating resources into other natural languages to share domain-specific cross-lingual knowledge. In this work, we conduct investigations on clinical text machine translation by examining multilingual neural network models using deep learning methods such as Transformer-based structures. Furthermore, to address the issue of language resource imbalance, we also carry out experiments using a transfer learning methodology based on massive multilingual pre-trained language models (MMPLMs). The experimental results on three sub-tasks including 1) clinical case (CC), 2) clinical terminology (CT), and 3) ontological concept (OC) show that our models achieved top-level performances in the ClinSpEn-2022 shared task on English-Spanish clinical domain data. Furthermore, our expert-based human evaluations demonstrate that the small-sized pre-trained language model (PLM) wins in the clinical domain fine-tuning over the other two extra-large language models by a large margin. This finding has never been previously reported in the field. Finally, the transfer learning method works well in our experimental setting using the WMT21fb model to accommodate a new Spanish language space that was not seen at the pretraining stage within WMT21fb itself – and this deserves further exploration for clinical knowledge transformation, e.g. investigation into more languages. These research findings can shed some light on domain-specific machine translation development, especially in clinical and healthcare fields. Further research projects can be carried out based on our work to improve healthcare text analytics and knowledge transformations. Our data will be openly available for research purposes at https://github.com/HECTA-UoM/ClinicalNMT
@@ 1 Introduction
In recent years, Healthcare Text Analytics (HECTA) have gained more attention from researchers across different disciplines, due to their impact on clinical treatment, decision-making, hospital operation, and their recently empowered capabilities. These developments have much to do with the latest development of powerful language models (LMs), advanced machine-learning (ML) technologies, and increasingly available digital healthcare data from the social media [Griciūtė et al., 2023, Oyebode et al., 2021, Luo et al., 2022] and discharged outpatient letters from hospital settings [Henry et al., 2020, Spasic and Nenadic, 2020, Percha, 2021].
@@
Intelligent healthcare systems have been deployed in some hospitals to support the clinicians’ diagnostics and decision-making regarding patients and their health problems [Noor et al., 2022, Qian et al., 2021]. Such usages include key information extraction (IE) from electronic health records (EHRs), normalisation to medical terminologies, knowledge graph (KG) construction, and relation extraction (RE) between symptoms (problems), diagnoses, treatments, and adverse drug events [Wu et al., 2022, Nguyen et al., 2023, Belkadi et al., 2023]. Some of these digital healthcare systems can also help patients self-diagnose in situations where no General Practitioners (GPs) and professional doctors are available [Wroge et al., 2018, Zhu et al., 2021].
@@
However, due to the language barriers and inequal accessibility of digital resources across languages, there is an urgent need for knowledge transfer, such as from one human language to another [Costa-jussà et al., 2022, Khoong and Rodriguez, 2022]. Thus, to help address digital health disparity, machine translation (MT) technologies can be of good use.
@@
MT is one of the earliest artificial intelligence (AI) branches dating back to the 1950s, and it has boomed in recent years along with other natural language processing (NLP) tasks due to the newly designed powerful Transformers learning model [Weaver, 1955, Vaswani et al., 2017, Devlin et al., 2018, Han et al., 2021a]. Several attention mechanisms designed in Transformer deep neural models have proven themselves capable of better learning from a large amount of available digital data compared to traditional statistical and neural network-based models [Han, 2022, Kuang et al., 2018, Han and Kuang, 2018].
@@
In this work, we investigate the state-of-the-art Transformer-based Neural MT (NMT) models in connection with clinical domain text translation, to facilitate digital healthcare and knowledge transfer with the workflow drawn in Figure 1. Being aware of some current development in the competition of language model sizes in the NLP field, we set up the following base models for comparison study: 1) a small-sized multilingual pre-trained Marian language model (s-MPLM), which was developed by researchers at the Adam Mickiewicz University in Poznan and by the NLP group at the University of Edinburgh [Junczys-Dowmunt et al., 2018a, Junczys-Dowmunt et al., 2018b]; and 2) a massive-sized multilingual pre-trained NLLB LM (MMPLM/xL-MPLM), developed by Meta-AI covering more than 200 languages [Costa-jussà et al., 2022]. In addition to this, we set up a third model to investigate the possibility of transfer learning in the clinical domain MT: 3) the WMT21fb model which is another MMPLM from Meta-AI but with a limited amount of pre-trained language pairs including from English to Czech, German, Hausa, Icelandic, Japanese, Russian, and Chinese, and the opposite [Tran et al., 2021].
@@
The testing language pairs of these translation models in our work are English $\leftrightarrow$ Spanish. As far as we know, there are no other language pairs of openly available resources in the clinical domain MT. We use the international shared task challenge data from ClinSpEn2022 “clinical domain Spanish-English MT 2022” for this purpose  [^1] https://codalab.lisn.upsaclay.fr/competitions/6696. ClinSpEn2022 was a sub-task of the BioMedical MT track at WMT2022 [Neves et al., 2022]. There are three translation tasks inside ClinSpEn2022 including i) clinical cases report; ii) clinical terms, and iii) ontological concepts from the biomedical domain.
@@
Regarding the evaluation of these LMs, we used the evaluation platform offered by the ClinSpEn2022 shared task including several automatic metrics such as BLEU, METEOR, ROUGE, COMET. However, the automatic evaluation results did not give any apparent differentiation between the models on some of the tasks. Furthermore, there are issues like inconsistency regarding model ranking across automatic metrics. To address these issues and give a high-quality evaluation, we performed an expert-based human evaluation on the three models using outputs of Task one “clinical case report”.
@@
Our experimental investigation shows that 1) the extra-large MMPLM does not necessarily win over the small-sized MPLM on clinical domain MT via fine-tuning; 2) our transfer-learning model works successfully for clinical domain MT task on language pairs that were not pre-trained for, but added with fine-tuning. The first finding can shed some light on the idea that in clinical domain-specific MT, it is better to do more data cleaning and fine-tuning rather than build extra large LMs. Our second finding tells us the capability of MMPLMs in generating a new language pair knowledge space for translating clinical domain text even though this language pair was unseen in the pre-training stage with our experimental settings. This can be useful to low-resource NLP, such as the work by [Almansor and Al-Ani, 2018, Islam et al., 2021].  [^2] This paper reports systematic investigation findings based upon the preliminary work from [Han et al., 2022a, Han et al., 2022b, Han et al., 2023].
@@
The rest of this article is organised as below: Section 2 surveys other works related to ours, including clinical domain MT and NLP, large LMs, and transfer learning. Section 3 details the three LMs we deployed for comparison study. Section 4 introduces the experimental work we carried out and automatic evaluation outcomes. Section 5 follows up with expert-based human evaluation and the results. Finally, Section 6 concludes our work with discussion.
@@ 2 Related Work
Applying NLP models to clinical healthcare has attracted much attention of researchers, such as the work on disease status prediction using discharge summaries by ?), temporal expressions and events extraction from clinical narratives using combined methods of rules and machine learning by ?) and recent deep-learning models by [Tu et al., 2023], Temporal Relation modelling on treatments using prompt engineering on GPT models by [Cui et al., 2023], using knowledge-based and data-driven methods for de-identification task in clinical narratives by ?), systematic reviews on clinical text mining and healthcare by ?) and ?), etc.
@@
However, using MT to help translate clinical text for knowledge transfer and improved clinical decision-making is still a relative novelty [Khoong and Rodriguez, 2022], even though it has proven its usefulness for assisting health communication especially with post-editing strategies [Dew et al., 2018]. This is partially the result of the sensitive nature of domain and high risk in clinical settings [Randhawa et al., 2013]. Some of the recent progress on using MT for clinical texts includes the work by ?) which leverages SNOMED-CT terms [Donnelly, 2006] and relations for MT between Basque and Spanish languages; ?) which applies NMT model to identify semantic concepts in “abundant interchangeable words” in clinical domain and their experimental result shows that NMT model can greatly improve the efficiency on extracting UMLS [Bodenreider, 2004] concepts from a single document by using 30 milliseconds in comparison to traditional regular expression-based methods which take 3 seconds; and ?) which uses NMT to simplify the typical multi-stage workflow on clinical report dictation and even correct the errors from speech recognition.
@@
With the prevalence of multilingual PLMs (MPLMs) developed from NLP fields, it becomes necessary to test their performances in the clinical domain of NMT. MPLMs have been adopted for many NLP tasks since the first emergence of the Transformer-based learning structure [Vaswani et al., 2017]. Among these, Marian is a small-sized MPLM led by Microsoft Translator based upon Nenatus NMT [Sennrich et al., 2017] with around 7.6 million parameters [Junczys-Dowmunt et al., 2018a]. At the same time, different research and development teams have been competing in recent years in terms of the size of their LMs such as the massive MPLMs (MMPLMs) WMT21fb and NLLB by Meta-AI that have the number of parameters set at 4.7 billion and 54 billion respectively [Tran et al., 2021, Costa-jussà et al., 2022]. To investigate the performances of these different models with varied model sizes towards clinical domain NMT with fine-tuning, we set up all three of these as our base models. To the best of our knowledge, our work is the first to compare small-size and extra-large MPLMs in the clinical domain of NMT.
@@
Close to the clinical domain, there is a biomedical domain MT challengethat has been organised along with the Annual Conference of MT (WMT) since 2016 [Bojar et al., 2016, Yeganova et al., 2021]. The historical biomedical MT tasks have covered corpus of biomedical terminologies, scientific abstracts from Medline, summaries of proposals for animal experiments, etc. In 2022, it was the first time that this Biomedical-MT shared task introduced clinical domain data for Spanish-English language pairs [Neves et al., 2022].
@@
As the WMT21fb model does not include Spanish in its pre-training, we also examined the transfer learning technology into the clinical domain NMT towards Spanish-English using the WMT21fb model. Transfer-learning [Alyafeai et al., 2020] has proved useful for text classification and relation extraction [Pomares-Quimbaya et al., 2021, Peng et al., 2019], and low-resource MT [Jiang et al., 2022] fields. However, to the best of our knowledge, we are the first to test clinical domain NMT via transfer learning using MMPLMs.
@@ 3 Experimental Designs
In this section, we introduce more information about the three MPLMs that we investigate in this work, i.e., Marian [Junczys-Dowmunt et al., 2018a], WMT21fb [Tran et al., 2021], and NLLB [Costa-jussà et al., 2022].
@@
Figure 3: Several Attention-based Transformer NMT structure [Vaswani et al., 2017].
@@ 3.1 Multilingual Marian NMT
First, we draw a training diagram of the original Marian model on its pre-training steps in Figure 2 according to [Junczys-Dowmunt et al., 2018a]. The pre-processing step includes tokenisation, true-casing, and Byte-Pair Encoding (BPE) for sub-words. The shallow training is to teach a mid-phase translation model to produce temporary target outputs for back-translation. Then, the back-translation step produces the same amount of input source sentences to enlarge the corpus. The deep-training step first uses four left-to-right models which can be RNN [Sennrich et al., 2017] or Transformer [Vaswani et al., 2017] structures, which is followed by four right-to-left models in the opposite direction. The ensemble-decoding step will generate the n-best hypothesis translations for each source input segment, which will be re-ranked using a re-scoring mechanism. Finally, in Marian NMT, there is an automatic post-editing step taken before the final output is produced. This step is also based on an end-to-end neural structure by modelling the set(MT-output, source sentence)$\rightarrow$“post-edited output” as introduced by ?).
@@ 3.2 Extra-Large Multilingual WMT21fb and NLLB
Instead of the optional RNN structure used in the Marian model, both WMT21fb and NLLB massive-sized multilingual PLMs (MMPLMs) adopted Transformer as the main methodology. As shown in Figure 3, Transformer’s main components for encoder include position encoding, Multi-Head Attention, and Feed-Forward Network with layer normalisation at both two steps. The decoder uses Masked Multi-Head Attention to constrain the generation model only taking into account the already generated text.
To increase the model capacity without making the extra-large model too slow for training, the WMT21fb model included “Sparsely Gated Mixture-of- Expert (MoE)” models, inspired by the work from ?), into the FFN layer of the Transformer, as shown in Figure 4. The MoE model will only pass a sub-set of model parameters into the next level, thus decreasing the computational cost. However, this dropout is done in a random manner.
@@
To further improve on this, the NLLB model designed a Conditional MoE Routing layer inspired by ?) to ask the MoE model to decide which tokens to dropout based on how computationally intensive/resource-heavy they are to process or based on their routing efficiency. This is achieved by a binary gate, which assigns weights to dense FNN $FFN_{shared}$ or MoE Gating, as in Figure 5. The Conditional MoE also removes language-specific parameters for learning.
In summary, the WMT21fb and NLLB models share very similar learning structures, the biggest difference being that WMT21fb used language-specific constrained learning. The WMT21fb model we applied is ‘wmt21-dense-24-wide.En-X’ (and X-En direction) which has 4.7 billion parameters  [^4] https://github.com/facebookresearch/fairseq/tree/main/examples/wmt21 and contains the language pairs English $\leftrightarrow$ Chinese, Czech, German, Hausa, Icelandic, Japanese, and Russian. The full NLLB model includes 200+ languages and has 54.5 billion parameters. Due to the computational restriction, we applied the distilled model of NLLB, i.e. NLLB-distilled, which has 1.3 billion parameters.
The WMT21fb model does not have Spanish among its trained language pairs, while NLLB includes Spanish as a high-resource language. This is a perfect setting for us to examine the transfer-learning technology on the clinical domain NMT by fine-tuning a translation model for the Spanish language on the WMT21fb model and comparing the output with the NLLB model (Spanish version).